OpenStudy (anonymous):

What is the greatest common factor of 52x5 and 12x4?

OpenStudy (anonymous):

52*5 and 12*4 52 = 13*4 12 = 4*3 52*5 = 13*4*5 12*4 = 3*4*4 gcf = 4

OpenStudy (anonymous):

i wont be 4x^5 or 4x^4 @sangya21

OpenStudy (anonymous):

The highest number that divides exactly into two or more numbers. It is the "greatest" thing for simplifying fractions! so 4 is common in both as for x^4 or x^5, there isnt any x mentioned in question